OpenStudy (anonymous):

Verify the following identity: sec(x)-sin(x)tan(x) = cos(x) Please include steps! :/ Thank you!

OpenStudy (anonymous):

Use tan x = sin/cos x, add and multiply through, then use cos^ x + sin^2 x =1

OpenStudy (anonymous):

tan(x) = sin(x)/cos(x) and sec(x) = 1/cos(x) substitution on original expression gives (1/cos(x)) - (sin(x) * (sin(x)/cos(x))) = cos(x) => 1/cos(x) - sin²(x)/cos(x) = cos(x) => (1 - sin²(x))/cos(x) = cos(x) => 1 - sin²(x) = cos²(x) sin²(x) + cos²(x) = 1 is an identity